Compute all the minors is and cofactors of (123 201 234)


1
Expert's answer
2021-05-24T15:21:03-0400

Given matrix is-


"A=\begin{bmatrix} 1&2&3\\2&0&1\\2&3&4\end{bmatrix}"



Minors are

"M_{11}=\begin{vmatrix} 0&1\\3&4\end{vmatrix}=-3\\[9pt]M_{12}=\begin{vmatrix} 2&1\\2&4\end{vmatrix}=6\\[9pt]M_{13}=\begin{vmatrix} 2&0\\2&3\end{vmatrix}=6\\[9pt]M_{21}=\begin{vmatrix} 2&3\\3&4\end{vmatrix}=-1\\[9pt]M_{22}=\begin{vmatrix} 1&3\\2&4\end{vmatrix}=-2\\[9pt]M_{23}=\begin{vmatrix} 1&2\\2&3\end{vmatrix}=-1\\[9pt]M_{31}=\begin{vmatrix} 2&3\\0&1\end{vmatrix}=2\\[9pt]M_{32}=\begin{vmatrix} 1&3\\2&1\end{vmatrix}=-5\\[9pt]M_{33}=\begin{vmatrix} 1&2\\2&0\end{vmatrix}=-4\\"



Cofactor is-


"C_{ij}=(-1)^{i+j}M_{ij}"


="\begin{bmatrix} -3&-6&6\\1&-2&1\\2&5&-4\end{bmatrix}"